OpenStudy (highschoolmom2010):

algebra 2 solving equations....i think i did it right

OpenStudy (highschoolmom2010):

\[15-g=23-2g\] \[15-g-15=23-2g-15\] -g=8-2g -g+2g=8 g=8

OpenStudy (oleg3321):

its correct

OpenStudy (oleg3321):

medal and fan please

OpenStudy (phi):

you can always check your answer. replace g with 8 in the original problem 15-8 = 23 - 2*8 7 = 23 - 16 7=7 which is true. That means g=8 is the correct answer.